# Diffusion Models Are Promising for Ab Initio Structure Solutions from Nanocrystalline Powder Diffraction Data

### v1

@@ Abstract
A major challenge in materials science is the determination of the structure of nanometer sized objects. Here we present a novel approach that uses a generative machine learning model based on a Diffusion model that is trained on 45,229 known structures. The model factors both the measured diffraction pattern as well as relevant statistical priors on the unit cell of atomic cluster structures. Conditioned only on the chemical formula and the information-scarce finite-size broadened powder diffraction pattern, we find that our model, PXRDnet, can successfully solve simulated nanocrystals as small as 10 Å across 200 materials of varying symmetry and complexity, including structures from all seven crystal systems. We show that our model can determine structural solutions with up to $81.5\%$ accuracy, as measured by structural correlation. Furthermore, PXRDnet is capable of solving structures from noisy diffraction patterns gathered in real-world experiments. We suggest that data driven approaches, bootstrapped from theoretical simulation, will ultimately provide a path towards determining the structure of previously unsolved nano-materials.
@@ Introduction
In practice, for sufficiently simple structures it is possible to solve a structure from powder diffraction data alone [3, 4]. In the case of nano-materials it is rarely possible to solve the structure ab initio [5, 6].
@@
We use as input a powder x-ray diffraction (PXRD) pattern, measured over a limited range of momentum transfer, $Q$, with peak broadening due to crystallite nano-sizing. We note that, once demonstrated, the approach, can also be applied to powder patterns that have no finite-size broadening, resulting in a broadly applicable method for structure solution from powder and nano-powder diffraction data. The method shows great promise despite the highly degraded information content in the input patterns.
@@
Although there are a number of works applying ML to problems in structure solution, we are not aware of any similar to our approach in method or breadth of applicability. [7, 8]. Much of the early work in this area used machine learning to predict isolated properties such as unit cell parameters, magnetism, phase composition and symmetry group of structures given PXRD patterns [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. We also note that our problem is distinct from that of structure prediction, where generative artificial intelligence (AI) techniques have been used successfully to generate novel (never seen before) material structures that are presumably energetically stable [22, 23, 24, 25], e.g., GNoME [26], UniMat [27], CDVAE [28], MatterGen [29].
ML has also been used to help with sub-tasks in the overall structure solution problem, such as structure solution for a limited number of of crystal systems (CrystalNet [30]) and close-packed structures (DeepStruc [31]), AI accelerated mining of existing structural databases (MLstructureMining [32]), or single-crystal structural solutions (Crysformer [33, 34]).
@@
We find that our use of diffusion models for a fully end-to-end structure solution tool is highly promising. It was able to find plausible structures from information degraded PXRD patterns across broad classes of materials from all seven crystal systems. It was designed to output the lattice parameters and fractional atomic coordinates of atom in in a unit cell of the structure, in a crystallographic setting that can be ingested by other programs that take crystal structures as inputs.
@@
Owing to the inherent uncertainty when presented with information-scarce diffraction patterns, PXRDnet, via its Langevin dynamics-inspired conditional generative process [35], generates multiple candidate structures that adhere to the input information. These candidates can be refined to search for global minimum structure solutions. It successfully found structure solutions for simulated nanocrystals as small as 10 Å, is successful across all seven crystal systems, and was shown to work on real-world experimentally gathered PXRD patterns.
@@
Data PDF Match $\mathbf{R_{wp}^{2}}$ Number
Source Correlation ($\uparrow$) Rate ($\uparrow$) ($\downarrow$) Materials
PXRDnet
MP-20 (10 Å) $0.798\pm 0.11$ $0.500$ $0.377\pm 0.26$ $200$
MP-20 (100 Å) $0.791\pm 0.12$ $0.455$ $0.364\pm 0.27$ $200$
Experimental $0.815\pm 0.09$ $0.400$ $0.460\pm 0.39$ $15$
Baseline: Latent Space Search
MP-20 (10 Å) $0.696\pm 0.10$ $0.180$ $0.643\pm 0.30$ $200$
MP-20 (100 Å) $0.701\pm 0.10$ $0.205$ $0.635\pm 0.30$ $200$
Experimental $0.693\pm 0.10$ $0.133$ $0.694\pm 0.26$ $15$
Baseline: Semi-Random
MP-20 (10 Å) $0.567\pm 0.15$ $0.005$ $0.986\pm 0.37$ $200$
MP-20 (100 Å) $0.570\pm 0.15$ $0.010$ $0.989\pm 0.38$ $200$
Experimental $0.580\pm 0.16$ $0.067$ $1.100\pm 0.37$ $15$
Table 1: Material Reconstruction Success Metrics: Data Source refers to the source of the data, where "MP-20 ($\{10,100\}$)" refers to PXRD patterns of simulated $\{10,100\}$-scale nanomaterials from the MP-20-PXRD dataset [36, 28], and "Experimental" refers to PXRD patterns from real-world experimental data. "Baseline (MP-20, Experimental)" refers to results from semi-random search baselines on MP-20 and the Experimental dataset, respectively. PDF Correlation refers to the Pearson correlation coefficient [37] between the simulated pair distribution functions of the true material and of the predicted material; it ranges from $0\rightarrow 1$, where higher is better; we calculate the mean and standard deviation over the testing set. Match Rate refers to the proportion of predicted structures that PyMatGen’s default material matching function [38] deemed to be a match with the ground truth; it ranges from $0\rightarrow 1$, where higher is better. $R_{wp}^{2}$ (otherwise known as the R-Factor) refers to the relative error between the PXRD pattern of the ground truth material and the PXRD pattern of the predicted material, where lower values are better [4]; we calculate the mean and standard deviation over the testing set.
Figure 1: Nanomaterial PXRD Patterns: We simulate nanoscale shrinkage via the sinc filter, thereby broadening the PXRD peaks (purple lines) from the ideal pattern (gray lines). To improve model performance, we create a smoother target PXRD pattern (dotted pink lines) that removes the sharp ripples at the expense of further broadening the diffraction pattern. This is done via an additional Gaussian filter after the sinc filter. The horizontal axis is $Q$-1), and the vertical axis is scaled intensity (where $1$ is maximal).
@@
Figure 2: Histogram (Distribution) of $\mathbf{R_{wp}^{2}}$: Corresponds to $R_{wp}^{2}$ results shown in Table 1. The horizontal axis is $R_{wp}^{2}$ (the R-factor), i.e., the relative error between the predicted material’s PXRD pattern and the ground truth material’s PXRD pattern. The vertical axis is the number of materials from the testing set that had $R_{wp}^{2}$ at a certain level. The green line with stars represents PXRDnet candidates, the blue line with circles represents latent space search candidates, and the red line with dots represents random baseline candidates. Panel A is from PXRD patterns from 10 Å nanomaterials; Panel B is from PXRD patterns from 100 Å nanomaterials.
@@ Results
Typically, a nanostructure is defined as that with crystalline size below 1000 Å [39]. To test the efficacy of our approach, we go two orders of magnitude below that and simulate PXRD patterns from crystalline sizes 10 Å and 100 Å with a mathematically principled filtering approach based on Fourier analysis [40] (described in Methods). As expected, the 10 Å  case has more peak broadening (and therefore, information degradation) [2, 3] than the 100 Å case. Visualizations of representative PXRD patterns are shown in Figure 1.
To our model, we input this nanoscale PXRD pattern and the chemical formula of materials from MP-20 (a subset of the Materials Project database) [28, 36].The model outputs a set of material structure predictions in the form of lattice parameters and fractional coordinates of atoms in the unit cell. We have trained separate models for each nanocrystalline size, for a total of two models. We also have two baseline models (which are really ablations on PXRDnet): (1) Latent Space Search, which searches and decodes a trained diffusion model’s latent space for materials that adhere to the PXRD pattern, without using any PXRD conditioning to generate the initial latent codes (2) Semi-Random, which creates a population of unit cells with random atomic positions and lattice parameters, and filters this population by adherence to the PXRD pattern. See Methods for more details.
See Figures 3, 4, and 5 for representative structure predictions and Table 1 for success metrics. In reviewing our results, we emphasize that the nanostructure problem currently has no end-to-end solution in the literature.
@@
Figures 3, 4, 5 show a representative sampling of PXRDnet outputs. We selected the solutions to put into Figures 3, 4, and 5 by uniformly sampling materials from PXRDnet’s testing set. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset. In this qualitative analysis we sorted these into ones that visually appear to be successes, partial successes and failures.
[…]
@@
In the field of structure solution from powder diffraction, a final structure refinement step is applied where a local search is carried out such that structural parameters (lattice parameters and fractional coordinates) are modified using non-linear regression methods to obtain the best fit to the measured powder diffraction pattern [44]. Finally, the chemical plausibility (e.g., interatomic distances, bond angles, missing or extra sites in organic compounds) of the resulting structure is tested using CIF-checkers, most notably IUCr’s CheckCIF [45]. These steps are not easily automated and are not high throughput, and so not appropriate for our purposes. They are also problematic for us because our ab initio approach does not first find the unit-cell with high accuracy as is done in the indexing process with high resolution powder diffraction methods. We tried the local refinement and found that the local search is very sensitive to having correct unit cell parameters, even when these are less relevant parameters for small nanoparticles that are our target here. We believe that these issues will be resolved in the future, but rather than address them here we turn to other measures of success that are high throughput. We note that these may not accurately reflect the actual degree of success of our method. They are more likely to result in false negatives than false positives and likely under-count the success rate, though this needs to be established. The metrics are described in the "Methods" section.
The quantitative results of Table 1 support the qualitative observations and show that PXRDnet is successful at reconstructing materials in MP-20. We compare our models against the aforementioned Latent Space Search and Semi-Random Search baselines (more details in Methods). PyMatGen has a structural matching algorithm [38] which we ran with the default parameters, as was the procedure in many prior works [28, 22, 29, 23, 24, 25]. This function classifies as structure as correct or not within some small angular and translational tolerances. To get the number in Table 1 we feed the top $5$ candidates from PXRDnet to the PyMatGen matching algorithm. If one of the structures is successful we deem this to be a successful structure solution by this measure.
We see that PXRDnet’s predictions are considerably better than the latent space search baseline, and better yet than the semi-random baseline. By this measure we can predict the correct structure almost half the time, whereas the latent space search baseline gets it right about one-fifth of the time, and the random search baseline gets virtually none of them right. The best match rate is $50.0\%$ on the 10 Å crystalline size. This is contrary to expectations, as conventional wisdom suggests that the smaller the crystal, the higher the peak broadening, and the lower the information content in the PXRD signal [2, 3]. While we do not wish to baselessly speculate on the reasons for this, we note that peak broadening makes for a smoother and less sparse optimization objective, which is better for machine learning, in our experience. At any rate, this is good news, because it means that the Diffusion model can make progress on a problem that was previously thought to be intractably difficult. For the 100 Å particles the results were lower but similar, sitting at $45.5\%$ match rate.
Taking a different approach to measuring success we can compare the atomic pair distribution function (PDF) of the candidate structures to that of the ground-truth using the Pearson correlation function [37] as the distance metric. Here 1 implies a perfect agreement and 0 that the two curves are uncorrelated. The PDF is useful as a 1D representation of the 3D structure and is a the Fourier transform of the powder diffraction pattern [4]. For each material, we report the single best PDF correlation out of PXRDnet’s top $5$ returned candidates.
Overall, the PDF correlations of the two PXRDnet models follow the same trend as the PyMatGen match rate. The model structures significantly out-perform both the latent space and semi-random baseline models, with the latent space baseline being somewhat closer to PXRDnet’s performance. When we compare the two PXRDnet models to each other, there is not that much difference, as they are all within one standard deviation of each other.
@@
Finally, we measure $R_{wp}^{2}$, the profile-weighted R-factor, which is the intensity-weighted sum of squares of the deviation of the predicted and ground truth powder patterns (see "Methods" for details). Here, lower values are better. This measure is used as a target function in Rietveld structure refinements, whence a highly successful fit typically has $R_{wp}^{2}$ below $10\%$ [3]. Just as with PDF correlation, for each material, we report the single best $R_{wp}^{2}$ out of PXRDnet’s top $5$ returned candidates.
@@
This measure should perform similarly to the Pearson measure and it does. The two PXRDnet models perform at essentially the same level, and they both considerably outperform both the baseline models. Before model refinement, even from conventional structure solution methods from well crystallized samples, this measure can be considerably higher than 10%, even for the correct structure. This shows that our method, while far from perfect, can give decent structural candidates for a considerable number of materials.
@@
Figure 2 shows the distribution of $R_{wp}^{2}$ for PXRDnet (green), the latent search baseline (blue), and the random search baseline (red). We see that for both PXRDnet models, the peaks of the distributions are concentrated around relatively low $R_{wp}^{2}$, many below 10%, with rightward tails representing higher $R_{wp}^{2}$. In contrast, the latent search baselines peak at higher $R_{wp}^{2}$, with very few of their $R_{wp}^{2}$ crossing the 10% threshold. Similarly to the PXRDnet distribution, the latent search distribution is also somewhat right-skewed. Finally, the random search baselines have the highest $R_{wp}^{2}$ of all, with very little skew. This highlights the advantages of diffusion models in finding plausible structural candidates, and further emphasizes the power of PXRDnet specifically in generating high-quality structural candidates.
@@
Figure 3: PXRDnet Successful Cases: The leftmost column is the ground truth crystal structure. The other columns show PXRDnet’s reconstructed crystal structures from simulated nanocrystal PXRD patterns of diameter 10, 100 Å. For convenient visualization of some examples, we have tessellated their unit cells into super cells. We selected the solutions to put into Figures 3, 4, and 5 by uniformly sampling materials from PXRDnet’s testing set. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset.
Figure 4: PXRDnet Semi-Successful Cases: The leftmost column is the ground truth crystal structure. The other columns show PXRDnet’s reconstructed crystal structures from simulated nanocrystal PXRD patterns of diameter 10, 100 Å. For convenient visualization of some examples, we have tessellated their unit cells into super cells. We selected the solutions to put into Figures 3, 4, and 5 by uniformly sampling materials from PXRDnet’s testing set. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset.
Figure 5: PXRDnet Unsuccessful Cases: The leftmost column is the ground truth crystal structure. The other columns show PXRDnet’s reconstructed crystal structures from simulated nanocrystal PXRD patterns of diameter 10, 100 Å. For convenient visualization of some examples, we have tessellated their unit cells into super cells. We selected the solutions to put into Figures 3, 4, and 5 by uniformly sampling materials from PXRDnet’s testing set. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset.
@@
Match Rate ($\uparrow$)
Cubic ($52$) $0.769$ $0.788$ $0.481$ $0.481$
Tetragonal ($31$) $0.516$ $0.419$ $0.097$ $0.258$
Orthorhombic ($35$) $0.571$ $0.371$ $0.086$ $0.057$
Trigonal ($23$) $0.391$ $0.348$ $0.130$ $0.000$
Hexagonal ($12$) $0.250$ $0.417$ $0.000$ $0.167$
Monoclinic ($37$) $0.297$ $0.270$ $0.054$ $0.108$
Triclinic ($10$) $0.100$ $0.100$ $0.000$ $0.000$
PDF Correlation ($\uparrow$)
Cubic ($52$) $0.834\pm 0.12$ $0.834\pm 0.15$ $0.630\pm 0.10$ $0.668\pm 0.12$
Tetragonal ($31$) $0.773\pm 0.08$ $0.762\pm 0.11$ $0.690\pm 0.09$ $0.681\pm 0.08$
Orthorhombic ($35$) $0.796\pm 0.10$ $0.790\pm 0.12$ $0.719\pm 0.09$ $0.717\pm 0.08$
Trigonal ($23$) $0.778\pm 0.11$ $0.781\pm 0.11$ $0.695\pm 0.08$ $0.681\pm 0.09$
Hexagonal ($12$) $0.725\pm 0.12$ $0.721\pm 0.10$ $0.695\pm 0.06$ $0.691\pm 0.07$
Monoclinic ($37$) $0.796\pm 0.10$ $0.774\pm 0.10$ $0.746\pm 0.09$ $0.741\pm 0.09$
Triclinic ($10$) $0.841\pm 0.04$ $0.828\pm 0.05$ $0.788\pm 0.07$ $0.793\pm 0.05$
@@
Cubic ($52$) $0.692\pm 0.54$ $0.600\pm 0.54$ $1.745\pm 0.41$ $1.674\pm 0.41$
Tetragonal ($31$) $0.997\pm 0.42$ $1.024\pm 0.55$ $1.434\pm 0.40$ $1.484\pm 0.47$
Orthorhombic ($35$) $0.820\pm 0.39$ $0.818\pm 0.48$ $1.176\pm 0.43$ $1.159\pm 0.49$
Trigonal ($23$) $0.880\pm 0.36$ $0.983\pm 0.52$ $1.386\pm 0.48$ $1.338\pm 0.36$
Hexagonal ($12$) $1.173\pm 0.33$ $1.238\pm 0.40$ $1.312\pm 0.30$ $1.355\pm 0.30$
Monoclinic ($37$) $0.720\pm 0.33$ $0.808\pm 0.51$ $1.205\pm 0.53$ $1.143\pm 0.55$
Triclinic ($10$) $1.034\pm 0.51$ $1.083\pm 0.55$ $1.215\pm 0.36$ $1.141\pm 0.40$
Table 2: Results by Crystal System: Crystal System (#) refers to the crystal system name, and the number of materials we have in the dataset of that system. Interpretation of metrics and method names remains the same as Table 1. All results shown are for MP-20. Because the semi-random baseline is so poor, we omit it, and only show the latent space search baseline with PXRDnet.
@@
Unsurprisingly, and in line with previous work [30], we find that the cubic system is the easiest to solve, with match rates up to $80.8\%$ (as compared to a sub-$50\%$ overall match rate). This trend holds, regardless of nanomaterial crystal size. We also find that different nanomaterial crystal sizes are better suited for different crystal systems. For instance, the 10 Å size does very well on the orthorhombic and the tetragonal systems (which both have $90^{\circ}$ lattice angles), predicting the correct material above $50\%$ of the time by this measure, while the baseline models do not reach this level of success. Then, the 100 Å crystal size gives good results for hexagonal. The success rate does decrease with decreasing symmetry of the structure, with the triclinic crystal system performing the least well according to the PyMatGen metric. However, we note that this trend is much less apparent when looking at the PDF correlation as a metric. The PDF is more sensitive to the relative positions of atoms than to crystal symmetry. In this sense the PDF may be a better measure of success for small nanocrystals and atomic clusters. For the PDF metric there is less clear of a trend from higher symmetry to lower symmetry ground-truth structure. Caution must be exercised here due to there being only $10$ triclinic materials in the testing set.
[…]
Figure 6: Experimental Data: The leftmost column is the experimentally observed PXRD pattern, where the horizontal axis is $Q(\text{\leavevmode\nobreak\ \AA}^{-1}$), and the vertical axis is the scaled intensity of the pattern. The middle column is the ground truth crystal structure. The rightmost column is the predicted crystal structure. For convenient visualization of some examples, we have tessellated the unit cells. The labels are the chemical formula, Pearson’s correlation [37] between ground truth and simulated pair distribution functions, whether or not PyMatGen [38] determined the predicted structure to match the ground truth, and $R_{wp}^{2}$.
Figure 7: Experimental Data (Cont.): The leftmost column is the experimentally observed PXRD pattern, where the horizontal axis is $Q(\text{\leavevmode\nobreak\ \AA}^{-1}$), and the vertical axis is the scaled intensity of the pattern. The middle column is the ground truth crystal structure. The rightmost column is the predicted crystal structure. For convenient visualization of some examples, we have tessellated the unit cells. The labels are the chemical formula, Pearson’s correlation [37] between ground truth and simulated pair distribution functions, whether or not PyMatGen [38] determined the predicted structure to match the ground truth, and $R_{wp}^{2}$.
Figure 8: Experimental Data (Cont.): The leftmost column is the experimentally observed PXRD pattern, where the horizontal axis is $Q(\text{\leavevmode\nobreak\ \AA}^{-1}$), and the vertical axis is the scaled intensity of the pattern. The middle column is the ground truth crystal structure. The rightmost column is the predicted crystal structure. For convenient visualization of some examples, we have tessellated the unit cells. The labels are the chemical formula, Pearson’s correlation [37] between ground truth and simulated pair distribution functions, whether or not PyMatGen [38] determined the predicted structure to match the ground truth, and $R_{wp}^{2}$.
@@
Figure 9: Rietveld Refinement Results: We conducted Rietveld refinement on $20$ materials. Panel A shows the results for 10 Å nanocrystal size; Panel B shows the results for 100 Å nanocrystal size. The horizontal axis is the calculated $R_{wp}^{2}$ before refinement; the vertical axis is the calculated $R_{wp}^{2}$ after refinement. One outlier (2.4,0.05)($\sim$ 2.4$\sim$ 0.05)( ∼ 2.4 , ∼ 0.05 ) is excluded from Panel B. For 10 Å nanocrystal size, pre-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.526\pm 0.367$, and post-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.418\pm 0.224$. For 100 Å nanocrystal size, pre-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.478\pm 0.490$ (including outlier), and post-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.119\pm 0.075$ (including outlier).
In light of the discussion above about model refinement as a final step in conventional structure solution, we attempted a Rietveld refinement on a subset of 20 candidates from PXRDnet.

### v2

# Ab Initio Structure Solutions from Nanocrystalline Powder Diffraction Data via Diffusion Models

@@ Abstract
A major challenge in materials science is the determination of the structure of nanometer sized objects. Here we present a novel approach that uses a generative machine learning model based on diffusion processes that is trained on 45,229 known structures. The model factors both the measured diffraction pattern as well as relevant statistical priors on the unit cell of atomic cluster structures. Conditioned only on the chemical formula and the information-scarce finite-size broadened powder diffraction pattern, we find that our model, PXRDnet, can successfully solve simulated nanocrystals as small as 10 Å across 200 materials of varying symmetry and complexity, including structures from all seven crystal systems. We show that our model can successfully and verifiably determine structural candidates four out of five times, with average error among these candidates being only 7% (as measured by post-Rietveld refinement R-factor). Furthermore, PXRDnet is capable of solving structures from noisy diffraction patterns gathered in real-world experiments. We suggest that data driven approaches, bootstrapped from theoretical simulation, will ultimately provide a path towards determining the structure of previously unsolved nano-materials.
@@ Introduction
In practice, for sufficiently simple structures it is possible to solve a structure from powder diffraction data alone [3, 4]. Unsurprisingly, in the past year, a few concurrent works have tackled this setting of the problem for non-nanoscale materials [5, 6, 7]. Yet, in the case of nano-materials it is rarely possible to solve the structure ab initio – we tackle this previously unsolved problem [8, 9].
@@
We use as input a powder x-ray diffraction (PXRD) pattern, measured over a limited range of momentum transfer, $Q$, with peak broadening due to crystallite nano-sizing. We note that, once demonstrated, the approach can also be applied to powder patterns that have no finite-size broadening, resulting in a broadly applicable method for structure solution from powder and nano-powder diffraction data. The method shows great promise despite the highly degraded information content in the input patterns.
@@
Although there are a number of works applying ML to problems in structure solution, we are not aware of any on par with our approach in breadth of applicability [10, 11]. Much of the early work in this area used machine learning to predict isolated properties such as unit cell parameters, magnetism, phase composition and symmetry group of structures given PXRD patterns [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. We also note that our problem is distinct from that of structure prediction, where generative artificial intelligence (AI) techniques have been used successfully to generate novel (never seen before) material structures that are presumably energetically stable [26, 27, 28, 29], e.g., GNoME [30], UniMat [31], CDVAE [32], MatterGen [33]. ML has also been used to help with sub-tasks in the overall structure solution problem, such as structure solution for a limited number of of crystal systems (CrystalNet [34]) and close-packed structures (DeepStruc [35]), AI accelerated mining of existing structural databases (MLstructureMining [36]), single-crystal structural solutions (Crysformer [37, 38]), or structure solution with single-element materials [39].
Furthermore, in the past year, a few concurrent works employ generative AI approaches to solve structures end-to-end from PXRD [6, 7, 5]. However, the crucial limitation of all these prior works is that they only solve PXRD patterns without finite-size effects. As previously noted, structure determination in this non-nanoscale setting is considered by materials scientists to be more or less a solved problem [3, 4]. On the other hand, structure determination from nanoscale materials is significantly harder from a physical perspective, due to peak broadening caused by the finite crystal size [40, 9, 8]. Another major shortcoming of these works is that their code is not (yet) publicly available, thereby limiting reproducibility and/or comparisons to their work.
@@
We find that our use of diffusion models for a fully end-to-end structure solution tool is highly promising. It was able to find plausible structures from information degraded PXRD patterns across broad classes of materials from all seven crystal systems. It was designed to output the lattice parameters and fractional atomic coordinates of atoms in a unit cell of the structure, in a crystallographic setting that can be ingested by other programs that take crystal structures as inputs.
@@
Owing to the inherent uncertainty when presented with information-scarce diffraction patterns, PXRDnet, via its Langevin dynamics-inspired conditional generative process [41], generates multiple candidate structures that adhere to the input information. These candidates can then be refined using local search methods, finding the global minimum structure solution if the model places the candidate anywhere within its basin of attraction. It successfully found structure solutions for simulated nanocrystals as small as 10 Å, is successful across all seven crystal systems, and was shown to work on real-world experimentally gathered PXRD patterns.
@@
Data Source PXRDnet CDVAE-Search Number
(Ours) (Baseline) Materials
MP-20 (10 Å) $\mathbf{0.325}\pm 0.283$ $0.647\pm 0.301$ $200$
MP-20 (100 Å) $\mathbf{0.324}\pm 0.261$ $0.591\pm 0.273$ $200$
Experimental $\mathbf{0.383}\pm 0.330$ $0.634\pm 0.311$ $15$
Table 1: $\mathbf{R_{wp}^{2}}$ (Material Reconstruction Success): $\mathbf{R_{wp}^{2}}$ (otherwise known as the R-Factor or Residuals Function) refers to the (pre-Rietveld refinement) relative error between the PXRD pattern of the ground truth material and the PXRD pattern of the predicted material, where lower values are better [4]; we calculate the mean and standard deviation over the testing set. Data Source refers to the source of the data, where "MP-20 ($\{10,100\}$)" refers to PXRD patterns of simulated $\{10,100\}$-scale nanomaterials from the MP-20-PXRD dataset [42, 32], and "Experimental" refers to PXRD patterns from real-world experimental data. “CDVAE-Search" refers to results from latent space exploration of the unconditional CDVAE diffusion model.
Figure 1: Nanomaterial PXRD Patterns: We simulate nanoscale shrinkage via the sinc2 filter, thereby broadening the PXRD peaks (purple lines) from the ideal pattern (gray lines). To improve model performance, we create a smoother target PXRD pattern (dotted pink lines) that removes the sharp ripples at the expense of further broadening the diffraction pattern. This is done via an additional Gaussian filter after the sinc filter. The horizontal axis is $Q$-1), and the vertical axis is scaled intensity (where $1$ is maximal).
@@
Figure 2: Histogram (Distribution) of $\mathbf{R_{wp}^{2}}$: Corresponds to $R_{wp}^{2}$ results shown in Table 1. The horizontal axis is $R_{wp}^{2}$ (the R-factor), i.e., the relative error between the predicted material’s PXRD pattern and the ground truth material’s PXRD pattern. The vertical axis is the number of materials from the testing set that had $R_{wp}^{2}$ at a certain level. The green line with stars represents PXRDnet candidates, and the red line with dots represents CDVAE-Search candidates. Panel A is from PXRD patterns from 10 Å nanomaterials; Panel B is from PXRD patterns from 100 Å nanomaterials.
@@ Results
Typically, a nanostructure is defined as that with crystalline size below 1000 Å [43]. To test the efficacy of our approach, we go two orders of magnitude below that and simulate PXRD patterns from crystalline sizes 10 Å and 100 Å with a mathematically principled filtering approach based on Fourier analysis [44, 45] (described in Methods). As expected, the 10 Å  case has more peak broadening (and therefore, information degradation) [2, 3] than the 100 Å case. Visualizations of representative PXRD patterns are shown in Figure 1.
To our model, we input this nanoscale PXRD pattern and the chemical formula of materials from MP-20 (a subset of the Materials Project database) [32, 42]. The model outputs a set of material structure predictions in the form of lattice parameters and fractional coordinates of atoms in the unit cell. We have trained separate models for each nanocrystalline size, for a total of two models. We also have CDVAE-Search [32] as our baseline model (which is, in some sense, an ablation on PXRDnet): this searches and decodes a trained diffusion model’s latent space for materials that adhere to the PXRD pattern, without using any PXRD conditioning to generate the initial latent codes. We emphasize that at the moment, competing generative AI works for structure solution do not make their source code public, nor do they handle finite-size effects, thereby making CDVAE-Search the closest comparison we have. See Methods for more details.
See Figure 3 for representative structure predictions, Figure 4 for the corresponding PXRD patterns, and Table 1 for success metrics. In reviewing our results, we emphasize that the nanostructure problem currently has no end-to-end solution in the literature.
@@
Figure 3 shows a representative sampling of post-Rietveld refinement [3] PXRDnet outputs. We display the post-Rietveld versions because it is standard procedure in crystallography to do a local minimization refinement step, such as Rietveld refinement in the case of powder data, on structure candidates [3, 46]. We selected the solutions to put into Figures 3 by uniformly sampling materials from PXRDnet’s testing set. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset. Among the refined candidates for each material, we display the one with the lowest post-refinement R-factor (in a real-world experimental setting, this is knowable, since we have the ground truth PXRD pattern) [3].
@@
That being said, the crystallographic community generally uses $R_{wp}^{2}$, the profile-weighted R-factor, as the success metric. This is the intensity-scaled sum of squares of the deviation of the predicted and ground truth powder patterns (see "Methods" for details). It is commonly used as a target function in Rietveld structure refinements. Lower values are better, with the minimum value being 00, and $R_{wp}^{2}$ below $10\%$ generally considered successful [3, 4]. For each material, we report the single best $R_{wp}^{2}$ out of PXRDnet’s top $10$ returned candidates; this is reasonable, as in real-world scenarios, the ground-truth PXRD is measurable.
@@
The quantitative results of Table 1 support the qualitative observations and show that PXRDnet is successful at reconstructing materials in MP-20. We see that PXRDnet’s predictions are considerably better than the CDVAE-Search baseline (more details in “Methods"). The two PXRDnet models perform at essentially the same level pre-Rietveld refinement, and they both considerably outperform the CDVAE-Search baseline. We note that before model refinement, even from conventional structure solution methods from well crystallized samples, the $R_{wp}^{2}$ measure can be considerably higher than 10%, even for the correct structure. This shows that our method, while far from perfect, can give decent structural candidates for a considerable number of materials.
@@
Figure 2 shows the distribution of $R_{wp}^{2}$ for PXRDnet (green), the latent search baseline (blue), and the random search baseline (red). We see that for both PXRDnet models, the peaks of the distributions are concentrated around relatively low $R_{wp}^{2}$, many below 10%, with rightward tails representing higher $R_{wp}^{2}$. In contrast, the CDVAE-Search baselines peak at higher $R_{wp}^{2}$, with very few of their $R_{wp}^{2}$ crossing the 10% threshold. Similarly to the PXRDnet distribution, the latent search distribution is also somewhat right-skewed. This highlights the advantages of PXRDnet specifically in generating high-quality structural candidates.
@@
Figure 3: PXRDnet Structure Predictions: The leftmost column is the ground truth crystal structure. The other columns show PXRDnet’s reconstructed crystal structures (after Rietveld refinement) from simulated nanocrystal PXRD patterns of diameter 10, 100 Å. For convenient visualization of some examples, we have tessellated their unit cells into super cells and translated the origin of the cif files. These structures have been uniformly selected from the testing dataset. Thus, the spread of success and failure across these figures is roughly reflective of success and failure across the whole testing dataset.
Figure 4: PXRD Comparisons: Comparison (without finite size effects) of the ground truth PXRD pattern, the raw PXRDnet prediction’s PXRD pattern, and the Rietveld-refined PXRD pattern. Corresponds to results in Figure 3.
@@
Cubic ($52$) $0.185\pm 0.281$ $0.224\pm 0.311$ $0.671\pm 0.339$ $0.650\pm 0.344$
Tetragonal ($31$) $0.443\pm 0.315$ $0.378\pm 0.218$ $0.685\pm 0.268$ $0.630\pm 0.243$
Orthorhombic ($35$) $0.301\pm 0.208$ $0.380\pm 0.263$ $0.636\pm 0.296$ $0.500\pm 0.211$
Trigonal ($23$) $0.388\pm 0.228$ $0.344\pm 0.256$ $0.678\pm 0.220$ $0.615\pm 0.218$
Hexagonal ($12$) $0.416\pm 0.155$ $0.366\pm 0.195$ $0.510\pm 0.192$ $0.542\pm 0.152$
Monoclinic ($37$) $0.318\pm 0.160$ $0.328\pm 0.226$ $0.607\pm 0.307$ $0.564\pm 0.272$
Triclinic ($10$) $0.547\pm 0.521$ $0.368\pm 0.104$ $0.679\pm 0.383$ $0.590\pm 0.263$
Table 2: Results by Crystal System: Crystal System (#) refers to the crystal system name, and the number of materials we have in the dataset of that system. Interpretation of metrics and method names remains the same as Table 1. All results shown are for MP-20.
@@
Unsurprisingly, and in line with previous work [34], we find that the cubic system is the easiest to solve. This trend holds, regardless of nanomaterial crystal size. The success does decrease with decreasing symmetry of the structure, with the triclinic and tetragonal crystal systems performing poorly.
@@
Figure 5: Rietveld Refinement Results: We conducted Rietveld refinement on ten promising candidates each for $20$ uniformly selected materials, making for a total of $200$ structures refined at each nanocrystal size. Panel A shows the results for 10 Å nanocrystal size; Panel B shows the results for 100 Å nanocrystal size. We display the results corresponding to the best post-refinement $R_{wp}^{2}$ (which can be calculated without knowledge of the ground truth structure). The horizontal axis is the calculated $R_{wp}^{2}$ before refinement; the vertical axis is the calculated $R_{wp}^{2}$ after refinement. For 10 Å nanocrystal size, the pre-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.561\pm 0.388$, and the post-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.132\pm 0.166$. For 100 Å nanocrystal size, the pre-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.663\pm 0.473$, and post-refinement $R_{wp}^{2}$ has $\mu\pm\sigma=0.068\pm 0.066$.
To further improve the results, we attempted a standard Rietveld refinement on a uniformly selected subset of 20 structures solved by PXRDnet (including those displayed in Figure 3). For each structure, we fed PXRDnet’s ten highest-ranked candidates. The results displayed are for the best candidate per material post-Rietveld refinement (again, this is knowable, since the ground truth PXRD pattern is observed).